OpenStudy (anonymous):

Find three solutions of the equation y = 7x – 5. (1 point) • (3, –26), (–1, –12), (3, 16) • (–3, –26), (–12, –1), (16, 3) • (–3, –26), (3, 16), (1, –12) • (–3, –26), (–1, –12), (3, 16) plx help as fast u can

pooja195 (pooja195):

same thing as the last one you just have to plug them in

OpenStudy (anonymous):

7 * -5? is that what i do

pooja195 (pooja195):

\[ -26 = 7(3) – 5\] \[ -12 = 7(-1) – 5\] \[ 16 = 7(3) – 5\]

OpenStudy (anonymous):

so B

pooja195 (pooja195):

no its not B -12 * 7 -5 is not -1

OpenStudy (anonymous):

C sorry bit this is all confusing

pooja195 (pooja195):

no now you're just guessing :O ok we need to plug in each number like i have done above we know its not A or B did u do the work for C?

OpenStudy (anonymous):

no

pooja195 (pooja195):

\[ -26 = 7(-3) – 5\] does this equal -26

OpenStudy (anonymous):

no

pooja195 (pooja195):

ok so that one is good :) because 7*-3=-21 - 5 = -26 so its okay \[ 16 = 7(3) – 5\]

pooja195 (pooja195):

7 x 3 = ?

OpenStudy (anonymous):

21

pooja195 (pooja195):

good now -5 ?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

so its D

pooja195 (pooja195):

Yep! Yay!!! you got it!

OpenStudy (anonymous):

Thanks!!!!!

pooja195 (pooja195):

you're welcome :)